Potassium acid oxalate K2C2O4.3H2C2O4.4H20 can be oxidized by MNO4- in acidic medium.Calculate the volume of 0.1M KMnO4 reacting in acid solution with one gram of acid oxalate.

Grade:11

1 Answers

Khushi
5 years ago
here n factor of MnO-4= 7 also, n factor of Mn2+= 2so required n factor = 5n factor of k2c2o4.3HC2o4.4H20=8nowmilliequivalent of KMnO4=milliequivalent of potassium acid oxalate(k2c2o4.3HC2o4.4H20)n factor* molarity *volume=w/E * 1000E= molecular weight/ n factorE=molecular weight of k2c2o4.3HC2o4.4H20 / 8E=505/8W=1 gram (given in the question)5*0.1*v= 1/505/8 *1000v=80000/2525v=31.6